# Scattering of non-relativistic finite-size particles and puffy dark matter direct detection

### v1

@@ 1 Introduction
Dark matter dominates the matter content of the universe Bertone:2016nfn ; Tian:2020tur , with evidence primarily coming from gravitational interaction, such as the galaxy rotation curves, the bullet cluster collision, and the observation of the cosmic microwave background (CMB) Ostriker:1973uit ; Randall:2008ppe . The dark matter particle candidates can have a mass in a vast range, spanning from $10^{-22}\rm eV$ to several tens of the solar mass Lin:2019uvt , with the most popular one being the Weakly Interacting Massive Particle (WIMP) because it can naturally provide the relic abundance and is accessible at collider experiments Buchmueller:2017qhf ; Jungman:1995df ; Lee:1977ua . Unfortunately, so far no unambiguous WIMP evidence has been observed in laboratory searches GAMBIT:2017zdo ; XENON:2024wpa ; XENON:2025vwd , which makes a light or ultralight dark matter candidate increasingly attractive. Since lighter dark matter particles give lower recoil energies in scattering off nucleus, they are much harder to detect in laboratory experiments and thus more sensitive instruments or advanced experimental techniques are needed.
@@
For direct detection of light dark matter, several approaches have been considered: dark matter–electron scattering Essig:2017kqs ; XENON:2024znc , the Migdal effect Knapen:2020aky , cosmic ray–boosted dark matter (CRDM) Wang:2021jic ; LZ:2025iaw , levitated particle detectors Kilian:2024fsg ; Cheng:2023loy , and varying the target material from noble gases to superconductors or semiconductors Liang:2022xbu ; Hochberg:2016ajh ; Liang:2024xcx ; Kurinsky:2019pgb . Some strategies for detecting ultralight dark matter have also been proposed An:2024wmc . Anyway, for direct detection of light dark matter, an accurate calculation of the dark matter–nucleus scattering cross section is crucial. Typically, the traditional scattering cross section is modeled as a constant prefactor multiplied by a momentum-transfer-dependent term and a form factor that accounts for the finite size of the target nucleus Acevedo:2021kly ; Hardy:2015boa ; Bhoonah:2020dzs ; Cappiello:2020lbk . However, in the non-relativistic regime, non-perturbative effects in particle scattering should be taken into account in order to obtain a realistic cross section Bollig:2021psb ; Bollig:2024ipe . A notable example is the inclusion of non-perturbative effects in point-like dark matter self-scattering to address the small-scale structure problems in astronomy Tulin:2013teo . Furthermore, at low velocities, attractive interactions between particles may lead to bound-state formation. Non-perturbative effects can also significantly enhance the annihilation cross section of low-energy point-like particles through the Sommerfeld enhancement mechanism Lee:1977ua ; Beneke:2022rjv . On the other hand, the finite-size effect of particles may also play a critical role in their scattering. When considering the scattering of finite-size dark matter particles off target nuclei, an additional form factor associated with the size of the dark matter must be included, which leads to a larger cross section exclusion region compared to point-like dark matter Wang:2023xgm . This remains true even in direct detection of semi-relativistic dark matter like the CRDM Wang:2023wrx . Actually, the particle size effect may directly influence its low-energy scattering, e.g., Ref. Xu:2020qjk discussed how the nuclear radius affects the dark matter-nucleus scattering cross section, Ref. Wang:2023xii showed that the finite-size dark matter self-scattering cross section can be categorized into Born, resonant, and classical regimes due to size effects, while Ref. Wang:2023wbw examined the influence of size effects on the Sommerfeld enhancement in the annihilation cross section of finite-size dark matter.
@@
Therefore, both non-perturbative and finite-size effects should be considered when analyzing scattering between non-relativistic, finite-size particles Digman:2019wdm . In this work, we will study the scattering between different finite-size particles and apply this size effect to the calculation of scattering cross section in the direct detection of puffy dark matter. First, we will re-examine the interaction potential between finite-size particles. Traditionally, such a potential is modeled as the point-like interaction potential multiplied by a form factor that captures the finite-size effect Chu:2018faw . Consequently, the scattering cross section is often given by the tree-level quantum field theory result for point particles multiplied by a form factor, namely, $d\sigma/d\Omega=\left(d\sigma/d\Omega\right)_{\rm point-like}F_{\rm size}(q)$, where $q$ is the momentum transfer Gelmini:2002ez . This result is equivalent to that obtained from quantum mechanical methods Wang:2021tjf . However, in the non-relativistic regime, a more precise treatment requires using quantum mechanics in coordinate space, integrating over the spatial extent of the finite-size target nucleus to obtain the correct dark matter–nucleus interaction potential. We will perform a state-of-the-art computation that incorporates the particle size effect accurately. Then by solving the Schrödinger equation, we will obtain a more realistic scattering cross section between finite-size particles, which includes non-perturbative effects in the non-relativistic regime Xu:2020qjk .
@@ 2 Scattering of non-relativistic finite-size particles
Theoretically, the Compton wavelength characterizes the quantum fluctuation scale of a particle at rest and is related to its positional uncertainty. If a particle’s size exceeds its Compton wavelength, its position can be localized, which is a necessary condition for the existence of a finite-size particle. On the other hand, the de Broglie wavelength describes the wave-like nature of a moving particle. When this wavelength is larger than the size of the target particle being probed, the moving (probing) particle perceives the target as a point-like particle. Conversely, if the de Broglie wavelength of the moving particle is smaller than the size of the target particle, the internal structure of the target can be resolved Colquhoun:2020adl .
@@
Similarly, for non-relativistic scattering between particles, we must consider the effective range of interaction (also called the force range), which is typically characterized by the inverse of the mediator mass, $1/m_{\phi}$. When the force range is much smaller than the incident particle’s de Broglie wavelength ( $1/m_{\phi}\ll 1/mv$), the interaction must be treated quantum mechanically, and the scattering is dominated by the s-wave. When the force range is comparable to the de Broglie wavelength ($1/m_{\phi}\sim 1/mv$), resonant or bound-state phenomena may occur during scattering Grabowska:2018lnd . When the force range is much larger than the de Broglie wavelength, the scattering corresponds to a long-range interaction and the wave nature of the incident particle no longer affects the spatial distribution of the force, in which the interaction can be treated semiclassically and the field-theoretic methods are typically used to calculate the scattering cross section.
@@
If a non-relativistic particle has finite size, then, as discussed in Ref. Wang:2023xii , the interaction potential between two such finite-size particles exhibits a sharp cutoff beyond their spatial extent due to the size effect. This short-range behavior is analogous to (albeit distinct from) nuclear forces (here we only consider elastic scattering). For traditional long-range potentials such as Coulomb or Yukawa interactions, the short-range size effect fundamentally modifies the calculation of the scattering cross section between finite-size particles. Therefore, when dealing with scattering between different finite-size particles, one must reconsider the key parameters used in classifying scattering cross sections, rather than directly applying the classification used for point-like non-relativistic particles.
@@
Here, we introduce some conventions for finite-size particles (also called puffy particles). Take finite-size dark matter particle as an example: a bound-state dark matter particle is composed of N point-like dark matter constituents, each with mass $m_{\chi}$, resulting in a total mass of $\rm Nm_{\chi}$  [^1] Here we neglect the binding energy and do not consider any dark strong interaction between the constituents. However, the scattering may be changed enormously by dark strong interaction in some parameter space Wang:2021tjf ..
@@ 2.1 The modified Yukawa potential
where $\rho_{1}(r_{1})$ and $\rho_{2}(r_{2})$ are the charge density functions, and $r$ is the distance between the two center points of the particle charges (in our calculation we choose the charge center point of one particle like a nucleus or a nucleon as the origin of coordinates). In general, the charge density function of a particle can take various forms, such as dipole, tophat, and Gaussian. In Ref. Chu:2018faw , the Hamiltonian between finite-size particles was computed in momentum space and found to be largely insensitive to the specific form of the charge density function. Similarly, Ref. Wang:2023xii also showed that the impact of the charge density profile on the potential is negligible. Therefore, in this work, we consider only spherically symmetric finite-size particles with a tophat-type charge density, given by
@@ 2.3 Partial wave approach
The details of solving the Schrödinger equation are presented in Appendix B of Ref. Wang:2023xii . Here we just briefly describe the calculation. The initial conditions like $\chi_{l}(x_{i})=1$ and $\chi^{\prime}_{l}(x_{i})=(l+1)/x_{i}$ are set at a point $x_{i}$ near the origin. This leads to an angular momentum term dominating the Schrödinger equation, which is then solved within the range $x_{i}\leq x\leq x_{m}$, with $x_{m}$ being the maximum value of $x$ used in the numerical analysis. With the condition of asymptotic solution Eq. (18) and $x=x_{m}$, we have
@@
Finally, by substituting the phase shifts into Eq. (19), we obtain the momentum-transfer scattering cross section. For non-relativistic self-scattering between point particles, the partial wave method reveals that the self-scattering parameter space is divided into Born, resonance, and classical regimes Tulin:2013teo . The non-perturbative dynamics in the resonance and classical regimes can address the small-scale structure anomalies in cosmology. Similarly, when applying the partial wave method to calculate the self-scattering cross section of finite-size dark matter, it is found that contributions from partial waves with $l>0$ are non-negligible, and the cross section obtained solely from the Born approximation does not represent the true value.
@@ 3 Size effect in puffy dark matter detection
The scattering cross section between a finite-size dark matter particle and a target nucleus is often assumed to be proportional to the square of the number of constituent particles comprising the dark matter bound state multiplied by the point particle–nucleus scattering cross section. However, this does not consider the volume effect of the finite-size dark matter particle. As can be seen from the modified potential in Eq.(6) and Fig.2, the size effect has a significant impact on the scattering. Therefore, we should start from the fundamental interaction and perform a volume integration for finite-size dark matter particle to obtain a correctly modified Yukawa potential, and then provide a more accurate scattering cross section for puffy dark matter detection.
@@ 3.1 Point dark matter-nucleus scattering
We fix the dark matter velocity as 300 km/s, the proton mass as $m_{N}=0.938\rm~GeV$, and the reduced mass is defined as $\mu=(m_{\chi}m_{N})/(m_{\chi}+m_{N})$. By performing a random scan over the parameter space $(m_{\phi},\alpha,m_{\chi})$, we obtain the results shown in Fig. 5. The right panel shows that for a large size-to-range ratio of the target nucleus, the scattering essentially lies within the Born approximation region. This is because in this regime the potential energy is nearly zero and much smaller than the kinetic energy, and consequently the scattering cross section is similar to the traditional point dark matter result. From the left and middle panels we see that for a small size-to-range ratio of the target nucleus, the scattering cross section classification resembles the result of point dark matter self-scattering, which is divided into Born, resonance, and classical regimes. Numerically, this arises from the competition between kinetic and potential energies of the scattering particles. In the s-wave case, the potential energy exceeds the kinetic energy, while as the partial wave number $l$ increases, the kinetic energy gradually dominates. However, unlike self-scattering, the competition between potential and kinetic energies here originates from the finite-size effect that modifies the Yukawa potential into a piecewise constant form rather than from the divergence behavior of the Yukawa potential itself. Even for a very small size-to-range ratio of the target nucleus, the size effect introduces a relatively large initial potential value in the piecewise potential. Furthermore, comparing the left and middle panels, we see that as the size-to-range ratio of the target nucleus increases, the parameter $a$ corresponding to the resonance peak decreases, until the resonance disappears entirely at a sufficiently large ratio. Similar results and analyses can also be found in Ref. Xu:2020qjk , which studied resonance scattering between point-like dark matter and baryons.
@@ 3.2 Puffy dark matter-nucleus scattering
We define $y=R_{\chi}m_{\phi}$ and the dimensionless parameter $bf(y)$ can be obtained by substituting the potential Eq.(14) into the Born approximation condition Eq.(10) (due to the complexity of the expression, its explicit form is not presented here). Then, we scan the parameter space $(m_{\phi},\alpha,m_{\chi})$ and obtain the results shown in Fig. 6. For a given value of $y$, we have $n=y/(m_{\phi}\times 6.2~\rm GeV^{-1})$. Fig. 6 displays the contour plots for the dimensionless parameters ($bf(y),a\sqrt{y}/f(y)$) with different scattering cross section values and different $R_{\chi}m_{\phi}$ values. For larger $R_{\chi}m_{\phi}$ values such as $R_{\chi}m_{\phi}=10^{3}$, the vertical axis values approach zero and thus are not plotted here. The dark matter velocity is again taken as 300 km/s. Similar to point dark matter–nucleus scattering, appropriate parameter choice allows for the classification of scattering cross sections. For scattering between particles of different sizes, here we choose the horizontal axis as $bf(y)$ since it shows the Born approximation condition with values smaller than 1. The reason for choosing $a\sqrt{y}/f(y)$ as the vertical axis is explained in Ref.Wang:2023xii .

### v2

@@ 1 Introduction
Dark matter dominates the matter content of the universe Bertone and Hooper (2018); Tian (2020), with evidence primarily coming from gravitational interaction, such as the galaxy rotation curves, the bullet cluster collision, and the observation of the cosmic microwave background (CMB) Ostriker and Peebles (1973); Randall et al. (2008). The dark matter particle candidates can have a mass in a vast range, spanning from $10^{-22}\rm eV$ to several tens of the solar mass Lin (2019), with the most popular one being the Weakly Interacting Massive Particle (WIMP) because it can naturally provide the relic abundance and is accessible at collider experiments Buchmueller et al. (2017); Jungman et al. (1996); Lee and Weinberg (1977). Unfortunately, so far no unambiguous WIMP evidence has been observed in laboratory searches Athron and others (2017); Aprile and others (2024, 2025b), which makes a light or ultralight dark matter candidate increasingly attractive. Since lighter dark matter particles give lower recoil energies in scattering off nucleus, they are much harder to detect in laboratory experiments and thus more sensitive instruments or advanced experimental techniques are needed.
@@
For direct detection of light dark matter, several approaches have been considered: dark matter–electron scattering Essig et al. (2017); Aprile and others (2025a), the Migdal effect Knapen et al. (2021), cosmic ray–boosted dark matter (CRDM) Wang et al. (2022a); Aalbers and others (2025), levitated particle detectors Kilian and others (2024); Cheng et al. (2025), and varying the target material from noble gases to superconductors or semiconductors Liang et al. (2022); Hochberg et al. (2016); Liang et al. (2025); Kurinsky et al. (2019). Some strategies for detecting ultralight dark matter have also been proposed An et al. (2025). Anyway, for direct detection of light dark matter, an accurate calculation of the dark matter–nucleus scattering cross section is crucial. Typically, the traditional scattering cross section is modeled as a constant prefactor multiplied by a momentum-transfer-dependent term and a form factor that accounts for the finite size of the target nucleus Acevedo et al. (2022); Hardy et al. (2015); Bhoonah et al. (2021); Cappiello et al. (2021). However, in the non-relativistic regime, non-perturbative effects in particle scattering should be taken into account in order to obtain a realistic cross section Bollig and Vogl (2022); Bollig (2024). A notable example is the inclusion of non-perturbative effects in point-like dark matter self-scattering to address the small-scale structure problems in astronomy Tulin et al. (2013). Furthermore, at low velocities, attractive interactions between particles may lead to bound-state formation. Non-perturbative effects can also significantly enhance the annihilation cross section of low-energy point-like particles through the Sommerfeld enhancement mechanism Lee and Weinberg (1977); Beneke et al. (2023). On the other hand, the finite-size effect of particles may also play a critical role in their scattering. When considering the scattering of finite-size dark matter particles off target nuclei, an additional form factor associated with the size of the dark matter must be included, which leads to a larger cross section exclusion region compared to point-like dark matter Wang et al. (2023c). This remains true even in direct detection of semi-relativistic dark matter like the CRDM Wang et al. (2023b). Actually, the particle size effect may directly influence its low-energy scattering, e.g., Ref. Xu and Farrar (2023) discussed how the nuclear radius affects the dark matter-nucleus scattering cross section, Ref. Wang et al. (2023a) showed that the finite-size dark matter self-scattering cross section can be categorized into Born, resonant, and classical regimes due to size effects, while Ref. Wang et al. (2024) examined the influence of size effects on the Sommerfeld enhancement in the annihilation cross section of finite-size dark matter.
@@
Therefore, both non-perturbative and finite-size effects should be considered when analyzing scattering between non-relativistic, finite-size particles Digman et al. (2019). In this work, we will study the scattering between different finite-size particles and apply this size effect to the calculation of scattering cross section in the direct detection of puffy dark matter. First, we will re-examine the interaction potential between finite-size particles. Traditionally, such a potential is modeled as the point-like interaction potential multiplied by a form factor that captures the finite-size effect Chu et al. (2020). Consequently, the scattering cross section is often given by the tree-level quantum field theory result for point particles multiplied by a form factor, namely, $d\sigma/d\Omega=\left(d\sigma/d\Omega\right)_{\rm point-like}F_{\rm size}(q)$, where $q$ is the momentum transfer Gelmini et al. (2002). This result is equivalent to that obtained from quantum mechanical methods Wang et al. (2022b). However, in the non-relativistic regime, a more precise treatment requires using quantum mechanics in coordinate space, integrating over the spatial extent of the finite-size target nucleus to obtain the correct dark matter–nucleus interaction potential. We will perform a state-of-the-art computation that incorporates the particle size effect accurately. Then by solving the Schrödinger equation, we will obtain a more realistic scattering cross section between finite-size particles, which includes non-perturbative effects in the non-relativistic regime Xu and Farrar (2023).
@@ 2 Scattering of non-relativistic finite-size particles
Theoretically, the Compton wavelength characterizes the quantum fluctuation scale of a particle at rest and is related to its positional uncertainty. If a particle’s size exceeds its Compton wavelength, its position can be localized, which is a necessary condition for the existence of a finite-size particle. On the other hand, the de Broglie wavelength describes the wave-like nature of a moving particle. When this wavelength is larger than the size of the target particle being probed, the moving (probing) particle perceives the target as a point-like particle. Conversely, if the de Broglie wavelength of the moving particle is smaller than the size of the target particle, the internal structure of the target can be resolved Colquhoun et al. (2021).
@@
Similarly, for non-relativistic scattering between particles, we must consider the effective range of interaction (also called the force range), which is typically characterized by the inverse of the mediator mass, $1/m_{\phi}$. When the force range is much smaller than the incident particle’s de Broglie wavelength ( $1/m_{\phi}\ll 1/mv$), the interaction must be treated quantum mechanically, and the scattering is dominated by the s-wave. When the force range is comparable to the de Broglie wavelength ($1/m_{\phi}\sim 1/mv$), resonant or bound-state phenomena may occur during scattering Grabowska et al. (2018). When the force range is much larger than the de Broglie wavelength, the scattering corresponds to a long-range interaction and the wave nature of the incident particle no longer affects the spatial distribution of the force, in which the interaction can be treated semiclassically and the field-theoretic methods are typically used to calculate the scattering cross section.
@@
If a non-relativistic particle has finite size, then, as discussed in Ref. Wang et al. (2023a), the interaction potential between two such finite-size particles exhibits a sharp cutoff beyond their spatial extent due to the size effect. This short-range behavior is analogous to (albeit distinct from) nuclear forces (here we only consider elastic scattering). For traditional long-range potentials such as Coulomb or Yukawa interactions, the short-range size effect fundamentally modifies the calculation of the scattering cross section between finite-size particles. Therefore, when dealing with scattering between different finite-size particles, one must reconsider the key parameters used in classifying scattering cross sections, rather than directly applying the classification used for point-like non-relativistic particles.
@@
Here, we introduce some conventions for finite-size particles (also called puffy particles). Take finite-size dark matter particle as an example: a bound-state dark matter particle is composed of N point-like dark matter constituents, each with mass $m_{\chi}$, resulting in a total mass of $\rm Nm_{\chi}$  [^1] Here we neglect the binding energy and do not consider any dark strong interaction between the constituents. However, the scattering may be changed enormously by dark strong interaction in some parameter space Wang et al. (2022b)..
@@ 2.1 The modified Yukawa potential
where $\rho_{1}(r_{1})$ and $\rho_{2}(r_{2})$ are the charge density functions, and $r$ is the distance between the two center points of the particle charges (in our calculation we choose the charge center point of one particle like a nucleus or a nucleon as the origin of coordinates). In general, the charge density function of a particle can take various forms, such as dipole, tophat, and Gaussian. In Ref. Chu et al. (2020), the Hamiltonian between finite-size particles was computed in momentum space and found to be largely insensitive to the specific form of the charge density function. Similarly, Ref. Wang et al. (2023a) also showed that the impact of the charge density profile on the potential is negligible. Therefore, in this work, we consider only spherically symmetric finite-size particles with a tophat-type charge density, given by
@@ 2.3 Partial wave approach
The details of solving the Schrödinger equation are presented in Appendix B of Ref. Wang et al. (2023a). Here we just briefly describe the calculation. The initial conditions like $\chi_{l}(x_{i})=1$ and $\chi^{\prime}_{l}(x_{i})=(l+1)/x_{i}$ are set at a point $x_{i}$ near the origin. This leads to an angular momentum term dominating the Schrödinger equation, which is then solved within the range $x_{i}\leq x\leq x_{m}$, with $x_{m}$ being the maximum value of $x$ used in the numerical analysis. With the condition of asymptotic solution Eq. (18) and $x=x_{m}$, we have
@@
Finally, by substituting the phase shifts into Eq. (19), we obtain the momentum-transfer scattering cross section. For non-relativistic self-scattering between point particles, the partial wave method reveals that the self-scattering parameter space is divided into Born, resonance, and classical regimes Tulin et al. (2013). The non-perturbative dynamics in the resonance and classical regimes can address the small-scale structure anomalies in cosmology. Similarly, when applying the partial wave method to calculate the self-scattering cross section of finite-size dark matter, it is found that contributions from partial waves with $l>0$ are non-negligible, and the cross section obtained solely from the Born approximation does not represent the true value.
@@ 3 Size effect in puffy dark matter detection
The scattering cross section between a finite-size dark matter particle and a target nucleus is often assumed to be proportional to the square of the number of constituent particles comprising the dark matter bound state multiplied by the point particle–nucleus scattering cross section. However, this does not consider the volume effect of the finite-size dark matter particle. As can be seen from the modified potential in Eq.(6) and Fig. 2, the size effect has a significant impact on the scattering. Therefore, we should start from the fundamental interaction and perform a volume integration for finite-size dark matter particle to obtain a correct modified Yukawa potential, and then provide a more accurate scattering cross section for puffy dark matter detection. We will first re-calculate the scattering cross section between a point dark matter particle and a finite-size nucleus. We reproduced the results in Ref. Xu and Farrar (2023), which showed that such a cross section can be divided into Bonn, resonance and classical regions. Then we will extend the study of this scattering to the case of two finite-size particles, i.e., a finite-size dark matter particle and a finite-size nucleus. Finally, we will provide the results for the nugget dark matter particle.
@@ 3.1 Point dark matter-nucleus scattering
We fix the dark matter velocity as 300 km/s, the proton mass as $m_{N}=0.938\rm~GeV$, and the reduced mass is defined as $\mu=(m_{\chi}m_{N})/(m_{\chi}+m_{N})$. By performing a random scan over the parameter space $(m_{\phi},\alpha,m_{\chi})$, we obtain the results shown in Fig. 5. The right panel shows that for a large size-to-range ratio of the target nucleus, the scattering essentially lies within the Born approximation region. This is because in this regime the potential energy is nearly zero and much smaller than the kinetic energy, and consequently the scattering cross section is similar to the traditional point dark matter result. From the left and middle panels we see that for a small size-to-range ratio of the target nucleus, the scattering cross section classification resembles the result of point dark matter self-scattering, which is divided into Born, resonance, and classical regimes. Numerically, this arises from the competition between kinetic and potential energies of the scattering particles. In the s-wave case, the potential energy exceeds the kinetic energy, while as the partial wave number $l$ increases, the kinetic energy gradually dominates. However, unlike self-scattering, the competition between potential and kinetic energies here originates from the finite-size effect that modifies the Yukawa potential into a piecewise constant form rather than from the divergence behavior of the Yukawa potential itself. Even for a very small size-to-range ratio of the target nucleus, the size effect introduces a relatively large initial potential value in the piecewise potential. Furthermore, comparing the left and middle panels, we see that as the size-to-range ratio of the target nucleus increases, the parameter $a$ corresponding to the resonance peak decreases, until the resonance disappears entirely at a sufficiently large ratio. Similar results and analyses can also be found in Ref. Xu and Farrar (2023), which studied resonance scattering between point-like dark matter and baryons.
@@ 3.2 Puffy dark matter-nucleus scattering
We define $y=R_{\chi}m_{\phi}$ and the dimensionless parameter $bf(y)$ can be obtained by substituting the potential Eq.(14) into the Born approximation condition Eq.(10) (due to the complexity of the expression, its explicit form is not presented here). Then, we scan the parameter space ($m_{\phi}$, $\alpha$, $M_{\chi}$) with $M_{\chi}$ defined by $M_{\chi}=Nm_{\chi}$ and obtain the results shown in Fig. 6. Since we neglect the internal structure of the dark matter particle here, we consider only its total mass as a single parameter and do not specify the exact number of its constitutes. For a given value of $y$, we have $n=y/(m_{\phi}\times 6.2~\rm GeV^{-1})$. Fig. 6 displays the contour plots for the dimensionless parameters ($bf(y),a\sqrt{y}/f(y)$) with different scattering cross section values and different $R_{\chi}m_{\phi}$ values. For larger $R_{\chi}m_{\phi}$ values such as $R_{\chi}m_{\phi}=10^{3}$, the vertical axis values approach zero and thus are not plotted here. The dark matter velocity is again taken as 300 km/s. Similar to point dark matter–nucleus scattering, appropriate parameter choice allows for the classification of scattering cross sections. For scattering between particles of different sizes, here we choose the horizontal axis as $bf(y)$ since it shows the Born approximation condition with values smaller than 1. The reason for choosing $a\sqrt{y}/f(y)$ as the vertical axis is explained in Ref.Wang et al. (2023a).